OpenStudy (anonymous):

Graph each pair of parametric equations. x=2t y=t+5, -2 is less than or equal to "t" is less than or equal to 3

OpenStudy (tkhunny):

You would do well to construct a table. I might try t = -2, -1, 0, 1, 2, 3 and see if I need any more detail around any of those. In this case, it is easy enough to see what the result might be. The parametric equations are easily changed to a regular x-y relationship. y=t+5 ==> t = y-5 x=2t = 2(y-5) = 2y - 10 y = ½x + 5 in Slope-Intercept Form. In other words, your table should show you a line, so you won't find anything more interesting.